OpenStudy (toxicsugar22):

caculate the value of of (oh-) from the given (H30+) and label the solution as acidic or basic. a. 10^-8 M ________*10_____^ M b. 10^-11 M ______*10^ M

OpenStudy (anonymous):

\(\sf pH= -log [H^+]\\pOH= 14-pH\)

OpenStudy (toxicsugar22):

can you also explain to me how i would type this in a calucator

OpenStudy (toxicsugar22):

I have TI 30 calculator

OpenStudy (anonymous):

\(\sf [OH^-] = 10^{-pOH}\)

OpenStudy (anonymous):

not sure about your calculator.. but type "-" then "log" , enter the values from above, then say okay. that will be your pH...

OpenStudy (toxicsugar22):

i dont know how to do 10^-8 on calculator

OpenStudy (toxicsugar22):

do you know

OpenStudy (anonymous):

can you take a pic of your calculator ? it's either "EXP" or "x10^n", depending on your calc.

OpenStudy (toxicsugar22):

is it -8

OpenStudy (toxicsugar22):

the anwere

OpenStudy (anonymous):

yes for pH, but you are looking for OH-, so subtract the value that you get from 14

OpenStudy (toxicsugar22):

6

OpenStudy (toxicsugar22):

right

OpenStudy (anonymous):

yes! that will be the pOH last step would be \(\sf 10^ 6\), this will be your OH

OpenStudy (toxicsugar22):

SO IT IS -8 *10^6 M

OpenStudy (toxicsugar22):

or 8 *10^ 6 M

OpenStudy (anonymous):

nope, 10^6 is the answer already. you used pH=8 to solve pOH= 14-8 = 6 then you use pOH to solve for OH

OpenStudy (toxicsugar22):

how would i do that

OpenStudy (toxicsugar22):

can you please explain

OpenStudy (toxicsugar22):

8 *10^ 6 M

OpenStudy (toxicsugar22):

so this is the anwer then for the question right

OpenStudy (toxicsugar22):

1000000

OpenStudy (anonymous):

okay i'll do it step by step for you. "calculate the value of of (oh-) from the given (H30+) " Unknown: [OH-] Known: [H+] Analysis/Steps: 1. \(\sf pH= -log [H^+]\) -> use this to calculate for the pH since [H+] is given. \(\sf pH= -log [H^+]= - log [10^{-8}]= 8\) , so remember pH = 8 2. \(\sf pOH= 14-pH\) -> use this to calculate for pOH \(\sf pOH= 14-pH= 14-8=6\), therefore pOH is 6 3. \(\sf [OH^-] = 10^{-pOH}\) use this to calculate for your unknown. \(\sf [OH^-] = 10^{-pOH}=10^{-6}\) therefore your answer will be \(\sf [OH^-] = 10^{-6}\)

OpenStudy (anonymous):

now follow this steps with the second problem

OpenStudy (anonymous):

\(\sf [OH^-] = 10^{-6}\ or\ 1 \times 10 ^{-6} M\)

OpenStudy (toxicsugar22):

now would it be acid or basic

OpenStudy (anonymous):

if pH > 7 , the solution will be basic if pH = 7 , it will be neutral if pH < 7, it will be acidic so look at the value of the pH , we calculated and you tell me if its acidic or basic

OpenStudy (anonymous):

are you sure? what's the value of our pH?

OpenStudy (toxicsugar22):

basic

OpenStudy (toxicsugar22):

PH is 8

OpenStudy (anonymous):

right (: can you try doing the second one? I'll check if it's right

OpenStudy (toxicsugar22):

is it 1* 10^-3

OpenStudy (toxicsugar22):

i tried on my own

OpenStudy (anonymous):

yes! glad you got it ^_^ so will it be acidic or basic?

OpenStudy (toxicsugar22):

basic as PH IS 11

OpenStudy (toxicsugar22):

RIGHT

OpenStudy (anonymous):

yEp! !!

OpenStudy (anonymous):

do you understand it now?

OpenStudy (toxicsugar22):

YES THANKS

OpenStudy (anonymous):

no prob! :)